OpenStudy (anonymous):

evaluate the integral a=1 b=2 f(x)= 3/t^4

OpenStudy (anonymous):

thing you mean f(t) = 3/t^4

OpenStudy (anonymous):

\[\int\limits_{1}^{2} \frac{3}{t^4} dt \]

OpenStudy (anonymous):

\[= 3 \int\limits_{1}^{2} t^{-4} dt\]

OpenStudy (anonymous):

\[= 3( \frac{t^{-3}}{-3} ) \]

OpenStudy (anonymous):

between the limits

OpenStudy (anonymous):

\[- [ \frac{1}{t^3}] \] between 2 and 1

OpenStudy (anonymous):

\[= - ( [ 1/8] - [1/4] ) = (1/8)\]

OpenStudy (anonymous):

1/8